U.UULLOTUIR 1. Calculate the density of methane, CH4, at 20°C with 0.95 atm of pressure. PM d = RT
Answer #1

Temperature = T= 20C= 20+273 = 293K

pressure = P = 0.95 atm

Molar mass of Methane gas = CH4 = M = 16.0 g/mole

Gas constant = R = 0.0821 L-atm/mol-k

d= PM/RT

d= 0.95 x 16.0 / (0.0821 x 293)

d= 0.6318 g/L

Density of methane = d = 0.632 g/L